- Emergency Help
- Computer Training for All Staff
- Blog Policy and Procedure
- Middle Managers Meeting Minutes
George A. Smathers Libraries
University of Florida
P.O. Box 117000
Gainesville, FL 32611-7000
Telephone: (352) 273-2800
Fax: (352) 392-2354
Information Technology Staff Meeting Notes
December 6, 2011
December 6, 2011
Present: Shruti Bagri, Thomas Bielicke, Laurie Brennan, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Michael Jay, Chris Nicolich, Amy Polk, Patricia Ruwell, Josh Spurgin, Mark Sullivan, Evan Wack
Guest: Rachel Schipper
Ongoing IT Updates
- Holiday Party - Thursday, December 7, noon-2PM, Smathers Room 1A.
- IT Update will most likely be postponed for tomorrow, as the Holiday Party folks need to get into 1A to decorate.
- Video equipment for West IT Shop - order underway, in Sonia's queue. Adobe software ordered, waiting on Software Licensing.
- Packaging ARES and Smart Board software. Testing Illiad. Took a crack at ArcGIS, but shelved.
- Migration - West Circulation scheduled for tomorrow. Acq & Licensing tentatively set for next week.
- Switch for West cart received.
- MSL cart received. Negotiating switch swap with Dell.
- AV - 2 new group study displays in MSL. Meeting with Mark McAlister about media site in room 1A.
- VDI project - waiting on Dell pricing.
- BPOP in MSL changed - boot times tested much better, in the same range as the other Libraries.
- PGP - takes forever to encrypt. Most likely to stick with BitLocker.
- DNS changer virus - 5 infections in the last 2 days.
- Kuali OLE - Chris is out. Circ policy docs under review. Chris to begin specification writing. Next meeting scheduled for February at Duke.
- Print Journal Inventory Tracker - Chris and Winston will be meeting with Lois Widmer to bring her up to speed.
- ASERL Gap Analysis - Tom and Shruti are working on it until Shruti leaves, but it will slow down as Shruti's last day is December 23 and Tom will soon be on vacation until after the first of the year.
- Circulation switch to "Library Guest" accounts - Bridges rolled out their portion. Roles have been assigned to test people. Auto groups have not been created to transfer to group policy. Bridges is anxious for us to use this designation, but we will need a period of training. Switch probably won't happen until the end of January.
- ARES - talking to Angela Mott and Scott Fagen at Atlas. What has been done is unclear. Need to escalate.
- Broadcast tool - ongoing.
- VDI Testing - access granted to the physical blade in CNS test lab. ESXi5 and vSphere Manager 5 to be installed this week.
- Server security - working with IT Security to add additional security to all of our servers.
- FCLA duties - transitioning from Matthew to Cynthia and Logan for certain IT related duties.
- ALADN listserv - in final move process.
- Departmental VPN - 'firstname.lastname@example.org/libit. If you use VPN, go ahead and start using this path.
- Room Reservation System - extended hours issue. FCLA is supposed to fix it tonight.
- ALF/Storage website update in process.
- LAC website in process.
- IIS/Web folder cleanup in process.
- Panama Canal Museum on hold. Have not heard back from them. Once we get a Web Developer hires, we can start on it.
- FCLA implementation of SobekCM.
- UFDC/Sobek web application update - in process.
- DLC-to-Hathi Trust format - can do once Digital Programmer is hired.
- LibEx - meeting in the future to discuss.